OpenStudy (anonymous):

Find the equation of the line tangent to the graph of the given function at the point with the indicated x-coordinate. f(x)= sqrt(x) + 1/sqrt(x)+5; x = 4 y= need help with calculus! I dont seem to be getting very far with this problem :/

OpenStudy (anonymous):

plug it in?

OpenStudy (mimi_x3):

Find the derivative first, then find m by substituting x=4 in. Then substitute m in y-y1 = m(x-x1)

OpenStudy (anonymous):

so find the slope at the given point by taking the derivitive. 1/(2sqrtx)-x^(-3/2)/2 plug in x=4